find the radius of the curvature at (acostheta^3,asintheta^3) on the curve x^2/3+y^2/3=a^2/3


1
Expert's answer
2021-12-22T13:11:02-0500

radius of the curvature:

"R=\frac{(1+(dy/dx)^2)^{3/2}}{|d^2y/dx^2|}"


"2x\/3+2yy'\/3=0"

"y'=-\frac{x}{y}"


"y''=-\frac{y-xy'}{y^2}"


at point "(acos \theta^3 ,asin \theta ^3):"

"y'=-cot\theta^3"


"y''=-\frac{sin\theta^3-cos^2\theta^3/sin\theta^3}{sin^2\theta ^3}"


"R=\frac{sin^2\theta ^3(1+cot^2\theta^3)^{3/2}}{sin\theta^3-cos^2\theta^3/sin\theta^3}"